OpenStudy (anonymous):

Find the mass of the plate D bounded by the given lines if the surface density of the plate p(x,y)[g/cm^2] is the known function: p(x,y)=(2x+5y)/(x^2+y^2), and the plate D is bounded by the following lines; x^2+y^2=9; x^2+y^2=16; x=0; y=0 (x>0, y>0) Make a sketch, show the plate configuration and present your answer with the corresponding dimension assuming that spatial variables x and y are in cms.

OpenStudy (anonymous):

Okay, Do you have a clue what the plate currently looks like?

OpenStudy (anonymous):

No what I wrote was all that was given

OpenStudy (anonymous):

Okay, Give me a sec to work this out.

OpenStudy (anonymous):

Okay, are you still there. I think I got it.

OpenStudy (anonymous):

yep sure am

OpenStudy (anonymous):

Okay, so your desired mass M would be:

OpenStudy (anonymous):

You would be integrating only a section of the bounded area

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Yes but where did the boundary conditions of 3 and 4 come from?

OpenStudy (anonymous):

Okay, if you sketch the circle, one being with a radius of 3 and the other of a radius of 4.

OpenStudy (anonymous):

x^2+y^2=9; x^2+y^2=16; first circle has a radius of 3 and the second a radius of 4

OpenStudy (anonymous):

Yes sorry just saw that as its taken from the given lines

OpenStudy (anonymous):

and since you only have all positive values, you would be integrating only up to to pi/2

OpenStudy (anonymous):

first quadrant of the circle correct?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

That's what this states x=0; y=0 (x>0, y>0)

OpenStudy (anonymous):

x greater than zero and y greater than zero

OpenStudy (anonymous):

Man, I didnt see that but its so obvious when you explain it. Thanks for that! So my sketch is just the 2 circles shown and then the area inbetween the circles is what we are finding. What the question asking when it says "with the corresponding dimension"?

OpenStudy (anonymous):

Yes, there are 2 circles, and it is between these two circles.

OpenStudy (anonymous):

Nice! got that bit, what about the last bit of the question when it refers to the "dimension"?

OpenStudy (anonymous):